# Agentic Visualization: Extracting Agent-based Design Patterns from Visualization Systems

### v1

@@ Abstract
*[regular] Examples: InsightsFeed [23] enables analysts to keep track of evolving results during progressive data analysis. In addition to the visual exploration of incrementally processed data, users can also monitor the progress of the background computation and see how far along it is. DataSite [25] also uses background processes to support the human analysts in visual data exploration and help them actively monitor the results of these processes using a feed timeline. In ProReveal [29], analysts can start the exploration and employ agents with different analytical support to guard the analysis process as a large dataset is progressively explored. It uses a web-based user interface where human analysts can explore the dataset, employ the guards, and continue exploring while Monitoring the guards’ progress. They are notified when a change occurs and the condition is modified.
@@
Pattern #11: As multiple agents perform various tasks throughout the analysis process, Consolidating the outcomes becomes essential to help both humans analysts and agents to build upon these findings. Based on the consolidated results, further exploration, summarization, or sensemaking of the data can be done. For example, after Consolidating the findings from Scouting the dataset, human analysts may wish to deepen their exploration in specific areas of interest. This is especially useful during sensemaking, where human analysts can interpret and communicate important findings from consolidated information.
@@
*[regular] Examples: InsightLens [39] uses an LLM-agent-based framework that automatically records and organizes insights during conversational data analysis with a human and an LLM agent. Based on the outcome of the analysis, the user can monitor and navigate the conversation using visualizations and receive guidance on further data exploration without interrupting their workflow. DataDirector [35] uses multiple agents to automate the creation of animated videos from data. LLM-based agents autonomously interpret the data, analyze it and further spawns design agent roles to create final data videos. In this case, tasks are broken down into sub-tasks at each analysis step, and the results from these tasks serve as input to the next step. In the final step, the main controller agent consolidates the decisions made by the agents to generate a data video for the human analyst.
@@
To demonstrate the practical application of our proposed AV patterns, we present an example scenario of a Medical Research Monitoring system that addresses the growing challenge of keeping pace with rapidly expanding medical literature. Consider a research team in a medical domain, where thousands of new publications emerge continuously across various sources that may be relevant to the field of study or a certain treatment or medication. The volume and velocity of this research make comprehensive manual monitoring impossible, yet staying current is critical. Traditionally, extensive manual efforts are required to identify key papers and extract significant information in order to visualize the literature landscape and synthesize the latest study results. This complex scenario leads to several interconnected challenges:
@@
Challenge 1: Query formulation and seeding. Significant domain expertise is required to create comprehensive search strategies that capture relevant publications without overwhelming noise. The Forager pattern directly addresses this challenge by deploying multiple autonomous agents that can explore different search strategies simultaneously. Each Forager agent can be configured with complementary expertise, and these agents can autonomously refine their search parameters based on initial results. Additionally, the Analyst pattern can assist researchers in formulating initial queries by analyzing existing literature collections to suggest relevant search terms and identify potential gaps in current search strategies.
@@
Challenge 2: Scale and filtering. The challenge of processing the overwhelming volume of publications that appear daily requires efficient pre-filtering mechanisms to identify potentially relevant studies while avoiding critical omissions that could impact research completeness. The Scouting coordination pattern provides a solution for handling the massive scale of incoming publications. Multiple scouting agents can be deployed to independently explore different publication streams, with each applying domain-specific filtering criteria while maintaining awareness of overall research priorities. This approach, combined with Monitoring, enables researchers to track the progress of filtering operations across multiple databases simultaneously. The Progress Indicator communication pattern keeps researchers informed about processing status without overwhelming them with intermediate results, allowing them to focus on high-priority publications while background agents handle the volume.
@@
Challenge 3: Concept extraction and validation. The task of automatically identifying and extracting key findings, methodologies, and outcomes from diverse publication formats while ensuring accuracy and completeness across varied writing styles and reporting standards presents significant complexity. The Analyst pattern excels at automated concept extraction, using natural language processing and domain-specific knowledge to identify key findings across diverse publication formats. These agents can generate structured summaries of extracted concepts and present them through the Insight Timeline communication pattern. Additionally, Chart Creator agents can generate domain-specific visualizations such as temporal trends, citation networks, and medication-specific charts that highlight key elements and emerging topics. The Provenance Log pattern ensures that all extracted concepts maintain links to their source publications, enabling researchers to verify accuracy and trace findings back to original sources for validation.
@@
Challenge 4: Research cluster identification. The task of discovering complex relationships between studies—including methodological similarities, conflicting results, and emerging research directions—requires sophisticated analysis of both explicit citations and implicit conceptual connections. The Swarming coordination pattern is particularly well-suited for discovering complex relationships between studies. Multiple agents can work collaboratively to identify clusters based on different similarity metrics—methodological approaches, research populations, outcome measures, or citation networks. Each agent in the swarm contributes its specialized perspective while coordinating with others to build an understanding of research relationships. The Insight Timeline pattern facilitates sharing of discovered clusters among team members, enabling collaborative validation and refinement of identified data patterns.
@@
Challenge 6: Temporal monitoring and adaptation. The ongoing need for continuously updating the monitoring system as new research emerges and research priorities evolve requires adaptive strategies that can respond to shifting focus areas and changing research methodologies. The Monitoring coordination pattern provides continuous oversight of the evolving research landscape. Monitoring agents track changes in publication patterns, emerging topics, and shifting research priorities, alerting researchers when significant developments occur. The Progress Indicator pattern communicates the evolution of research areas over time, while the Insight Timeline serves as a living record of research developments. This temporal awareness enables the system to adapt its focus areas dynamically, ensuring that monitoring strategies remain aligned with current research needs and emerging opportunities.
@@
Agentic visualization systems must navigate the relationship between computational autonomy and human analytical control. As Shneiderman [13] points out, human versus AI control is not a zero-sum game, but can be thought of as a 2D design space where human and computer control are independent axes. This framing reveals opportunities where both human and AI agents can retain high degrees of control.
As AI systems continue to evolve, we will most likely witness the rise of agents throughout the visual analysis process. Due to their significant computational power, agents can address issues with scalability, reducing the time from data to insights. LLM-based agents have already taken on collaborator roles in fields like programming, acting as pair programmers, code reviewers, and automated testers. A similar collaborative role could extend to visual analytics, where agents support human analysts by relieving them of cognitively demanding or tedious tasks.
Consider scenarios where AI agents proactively provide background and contextual information for insights, make suggestions for analytical directions, and provide careful provenance and source information for all decisions while leaving ultimate decision-making and sensemaking to the human analyst. An Analyst agent might autonomously identify statistical outliers in a dataset, gather relevant contextual information about those outliers, and present this analysis with full methodological transparency; a Chart Creator might suggest one or several charts suitable for representing the findings; and a Storyteller might suggest a comprehensive storyline for presenting the high-level insights. However, in the end, the human analyst retains control over interpreting significance and choosing next steps. With communication approaches such as Insight Timeline, agents and human analysts will be able to interact seamlessly through natural and visual language alike.
@@
The shift toward agentic visualization raises questions about analytical responsibility, interpretability, and the preservation of human values that have long defined our field. When an AI agent discovers a pattern that influences policy decisions, questions of responsibility become complex. How do we maintain transparency when agents make autonomous analytical choices? What decisions should agents be allowed to make? How do we ensure that increased automation enhances rather than undermines human analytical reasoning? These concerns extend beyond traditional AI ethics frameworks.
The patterns presented in this work may be helpful to address these issues. For example, Provenance Log can maintain comprehensive logs of data sources and transformations applied to the data by both agents and users. Such provenance trails could ensure transparency and reproducibility of the analysis process. Additionally, quality and uncertainty metrics could be employed [29] to indicate the quality of the results from agents. Still, ultimately, domain experts are required to validate the quality of the results.

### v2

@@ Abstract
*[regular] Examples: InsightsFeed [23] enables users to keep track of evolving results during progressive data analysis. In addition to the visual exploration of incrementally processed data, users can also monitor the progress of the background computation and see how far along it is. DataSite [25] also uses background processes to support the users in visual data exploration and help them actively monitor the results of these processes using a feed timeline. In ProReveal [29], users can start the exploration and employ agents with different analytical support to guard the analysis process as a large dataset is progressively explored. It uses a web-based user interface where users can explore the dataset, employ the guards, and continue exploring while Monitoring the guards’ progress. They are notified when a change occurs and the condition is modified.
@@
Pattern #11: As multiple agents perform various tasks throughout the analysis process, Consolidating the outcomes becomes essential to help users, such as evangelist, storyteller and agents to build upon these findings. Based on the consolidated results, further exploration, summarization, or sensemaking of the data can be done. For example, after Consolidating the findings from Scouting the dataset, users may wish to deepen their exploration in specific areas of interest. This is especially useful during sensemaking, where users can interpret and communicate important findings from consolidated information.
@@
*[regular] Examples: InsightLens [39] uses an LLM-agent-based framework that automatically records and organizes insights during conversational data analysis with a human and an LLM agent. Based on the outcome of the analysis, the user can monitor and navigate the conversation using visualizations and receive guidance on further data exploration without interrupting their workflow. DataDirector [35] uses multiple agents to automate the creation of animated videos from data. LLM-based agents autonomously interpret the data, analyze it and further spawns design agent roles to create final data videos. In this case, tasks are broken down into sub-tasks at each analysis step, and the results from these tasks serve as input to the next step. In the final step, the main controller agent consolidates the decisions made by the agents to generate a data video for the users.
@@
To demonstrate the practical application of our proposed AV patterns, we present a Medical Research Monitoring scenario that addresses the challenge of keeping pace with rapidly expanding medical literature. Our preliminary implementation efforts can be found at https://my-evidentia.vercel.app/. We also plan to deploy this tool and conduct a longitudinal study to evaluate its performance over time. However, this evaluation is beyond the scope of this work. Consider a research team in a medical domain, where thousands of new publications emerge continuously across various sources that may be relevant to the field of study or a certain treatment or medication. The volume and velocity of this research make comprehensive manual monitoring impossible, yet staying current is critical. Traditionally, identifying key papers and extracting meaningful information to visualize the literature landscape and synthesize findings requires extensive manual effort. In this scenario, the primary user roles include a research scientist, a technical analyst, and an evangelist. Not all roles from the broader data science framework [4] are present, reflecting the current structure of the organization. This complex scenario leads to several interconnected challenges:
@@
Challenge 1: Query formulation and seeding. Crafting effective search strategies that capture relevant publications requires domain expertise, often from research scientists. The Forager pattern can support researchers by deploying multiple autonomous agents that can explore different search strategies simultaneously. Each Forager agent can be configured with complementary expertise, and these agents can autonomously refine their search parameters based on initial results. Additionally, the Analyst pattern can assist researcher scientists by analyzing existing literature collections to suggest relevant search terms and identify potential gaps in current search strategies.
@@
Challenge 2: Scale and filtering. Research scientists and technical analysts need to manage the overwhelming volume of daily publications by applying efficient filtering techniques to identify relevant studies without missing critical information. The Scouting coordination pattern provides a solution for handling the massive scale of incoming publications. Multiple scouting agents can be deployed to independently explore different publication streams, with each applying domain-specific filtering criteria while maintaining awareness of overall research priorities. This approach, combined with Monitoring, enables researchers and analysts keep track of the progress of filtering operations across multiple databases simultaneously. The Progress Indicator communication pattern keeps researchers informed about processing status without overwhelming them with intermediate results, allowing them to focus on high-priority publications while background agents handle the volume.
@@
Challenge 3: Concept extraction and validation. For technical analysts, automatically identifying and extracting key findings, methodologies, and outcomes from diverse publication formats is a complex task that inherently involves close collaboration with scientists to ensure accuracy and uphold reporting standards. The Analyst pattern excels at automated concept extraction, using natural language processing and domain-specific knowledge to identify key findings across diverse publication formats. These agents can generate structured summaries of extracted concepts and present them through the Insight Timeline communication pattern. Additionally, Chart Creator agents can generate domain-specific visualizations such as temporal trends, citation networks, and medication-specific charts that highlight key elements and emerging topics. The Provenance Log pattern ensures that all extracted concepts maintain links to their source publications, enabling analysts and researchers to verify accuracy and trace findings back to original sources for validation.
@@
Challenge 4: Research cluster identification. The research scientists are responsible for discovering complex relationships between studies, including methodological similarities, conflicting results, and emerging research directions. This requires sophisticated analysis of both explicit citations and implicit conceptual connections. The Swarming coordination pattern is particularly well-suited for discovering complex relationships between studies. Multiple agents can work collaboratively to identify clusters based on different similarity metrics—methodological approaches, research populations, outcome measures, or citation networks. Each agent in the swarm contributes its specialized perspective while coordinating with others to build an understanding of research relationships. The Insight Timeline pattern facilitates sharing of discovered clusters among team members, enabling collaborative validation and refinement of identified data patterns.
@@
Challenge 6: Temporal monitoring and adaptation. The ongoing need for continuously updating the monitoring system as new research emerges and research priorities evolve requires adaptive strategies that can respond to shifting focus areas and changing research methodologies. The Monitoring coordination pattern provides continuous oversight of the evolving research landscape. Monitoring agents track changes in publication patterns, emerging topics, and shifting research priorities, alerting researchers and evangelists when significant developments occur. The Progress Indicator pattern communicates the evolution of research areas over time, while the Insight Timeline serves as a living record of research developments. This temporal awareness enables the system to adapt its focus areas dynamically, ensuring that monitoring strategies remain aligned with current research needs and emerging opportunities.
@@
Agentic visualization systems must navigate the relationship between computational autonomy and human analytical control. As Shneiderman [13] points out, human versus AI control is not a zero-sum game, but can be thought of as a 2D design space where human and computer control are independent axes. This framing reveals opportunities where both human and agents can retain high degrees of control.
We are currently witnessing the rise of agents throughout the visual analysis process, and this is expected to grow. With their computational capabilities, agents help address scalability challenges and reduce the time from data to insight. Through our extracted design patterns and use case scenario, multiple agents will play a collaborative role by supporting diverse human users by taking over cognitively demanding or repetitive tasks.
Several commercial tools, such as Microsoft Power BI and Tableau, have started integrating agents that support users from data processing to communication. While promising, the performance, utility and accuracy of these agents require further study to assess how effectively they amplify human capabilities.
Depending on the level of agency, the agents may only react in response to user input, such as in the case of Jupybara [40],Data Formulator [32], or proactively monitor user behavior to make suggestions, such as in the case of Ava [30] and InsightLens  [39]. In reactive systems, like Jupybara [40], a multi-agent system responds with an ordered plan to fulfill complex user queries. To ensure full transparency, it is essential to provide additional communication information such as Provenance Log and Progress Indicator, especially with large datasets where processing might take longer. In multi-agent systems, coordination patterns also play a critical role.
In proactive scenarios, agents may provide background and contextual information for insights, make suggestions for analytical directions, and provide careful provenance and source information for all decisions while leaving ultimate decision-making and sensemaking to the human analyst.
@@
The shift toward agentic visualization raises questions about analytical responsibility, interpretability, and the preservation of human values that have long defined our field. When an agent discovers a pattern that influences policy decisions, questions of responsibility become complex. How do we maintain transparency when agents make autonomous analytical choices? What decisions should agents be allowed to make? How do we ensure that increased automation enhances rather than undermines human analytical reasoning? These concerns extend beyond traditional AI ethics frameworks.
The patterns presented in this work may be helpful to address these issues. For example, Provenance Log can maintain comprehensive logs of data sources and transformations applied to the data by both agents and users to ensure transparency and reproducibility of the analysis process. Still, ultimately, domain experts are required to validate the quality of the results.